ans:
(a) 25
(b) 267
(c) 314.7
(d) 2.12


9 )A thermal storage battery consists of 10 kg of a salt solution that changes phase at 80°C and has the following properties: solid to liquid latent heat 80 W-hr/kg liquid 900 J/kg-K 350 J/kg-K specific heat as specific heat as solid = The battery container is superinsulated with a heat loss rate of 3 W (assume constant). En- gine coolant flows through the battery at a rate of 0.09 kg/sec, and has a temperature of 110°C when the engine is running. Coolant can be approximated as water. Calculate: How long it takes for the battery solution to reach 80°C after the engine is turned off. [hrl (a) (b) How long the battery solution remains at 80°C as it changes phase. [hr]
Heat Transfer in Engines (c) How long until the battery cools to an ambient temperature of 10°C. [hr (d) How long the battery can supply coolant at 80°C when the engine is started at a tem- perature of 20°C. Assume the battery solution starts as all liquid at 80°C, and the en- gine coolant enters at 20°C and leaves at 80°C. [min. Cemale Problem 5 has been inoperative in
Answer #1

9epresen their uouo maming Alualon Heat opA = 3 W m 10 kg = q003/g K Waler > HA = 350J A fiave, Hene So the -ho le kottes y kA-lc 3x3600 25 h Batay Aol for will 2emain at o c t2 dusalion Heat loae 3 2 h cs Scanned with CamScannerTim taken fo latter, Caol do un goc jo ts oQ2 E 2 2 8 fokem T= total lim e = 251 2 ((O CI 22 S =314. 3 S R 3147hs. cs Scanneda)ese the siluolions tas mo o ut lo lhe K. N-hA/K wales - 8o T= goC - 20 C H eat los = 3 W- Oec Heat Aupphedfhom battery toHeat nejactad phase chamge X l s W- 800 Aupply coolant for Liattery The Cam 800x 360o ls 22 074 A H 4 2 2 607 2 2 min Cs Scan